### v1

# FED: Fast and Efficient Dataset Deduplication Framework with GPU Acceleration

Abstract
Dataset deduplication plays a crucial role in enhancing data quality, ultimately improving training performance and efficiency of LLMs. A commonly used method for data deduplication is the MinHash LSH algorithm. Recently, NVIDIA introduced a GPU-based MinHash LSH deduplication method, but it remains suboptimal, leaving room for further improvement in processing efficiency. This paper proposes a GPU-accelerated deduplication framework FED that optimizes MinHash LSH for GPU clusters and leverages computationally efficient and partially reusable non-cryptographic hash functions. FED significantly outperforms the CPU-based deduplication tool included in SlimPajama by up to 58.3 times and the GPU-based deduplication tool included in NVIDIA NeMo Curator by up to 8.6 times when processing 1 million documents with a node of four GPUs. Deduplication of 1.2 trillion tokens is completed in just 5.1 hours in a four-node, 16-GPU environment. The related code is publicly available on GitHub (https://github.com/mcrl/FED).
FED: Fast and Efficient Dataset Deduplication Framework
with GPU Acceleration
Youngjun Son youngjun@aces.snu.ac.kr                        Chaewon Kim chaewon@aces.snu.ac.kr                        Jaejin Lee jaejin@snu.ac.kr
**footnotetext: These authors contributed equally to this workfootnotetext: Graduate school of Data Science, Seoul National Universityfootnotetext: Department of Computer Science, Seoul National University
1 Introduction
Pretrained language models (PLMs) Vaswani et al. (2017) have demonstrated exceptional performance, enabling unprecedented various application advancements. It is well known that the larger the dataset, the better the PLM’s performance Hoffmann et al. (2022).
High-quality datasets are crucial for LLM training because they directly affect the language models’ performance, generalizability, and robustness Gunasekar et al. (2023); Li et al. (2023). A critical aspect of ensuring data quality is the elimination of redundant data entries, a process known as dataset deduplication. It is natural to consider that duplicated data in the training dataset can make training an LLM more time-consuming and resource-intensive. It may also introduce bias into the model’s learning process towards near-duplicates in the dataset Lee et al. (2022).
Duplicate detection methods can be broadly categorized into two approaches: exact matching, which identifies identical text, and approximate matching, which identifies documents with similar content. Approximate matching is significantly more computationally expensive and time-consuming Albalak et al. (2024). One commonly used approximate matching method is MinHash LSH Indyk and Motwani (1998), a simplified version of MinHash Broder (1997). However, widely adopted implementations of MinHash LSH are too slow to handle large-scale datasets effectively.
Recently, NVIDIA introduced NeMo-Curator Jennings et al. , a data processing toolkit that includes a GPU-accelerated deduplication tool. While it is faster than traditional CPU-based implementations, processing datasets with trillions of tokens still takes over a day. This fact emphasizes the need for further innovation in deduplication techniques to address the ever-increasing scale of PLM datasets.
This paper presents a GPU-based deduplication framework called FED, which significantly outperforms the existing deduplication tools. By leveraging a simple yet efficient hashing function and optimizing GPU kernel operations and communication, we have developed a framework capable of deduplicating a dataset of 1.2 trillion tokens in just 5.1 hours. The contributions of this paper are summarized as follows:
  • GPU-accelerated deduplication. We present a new GPU-accelerated deduplication framework, called FED, optimized for computation and communication. We successfully deduplicated a dataset with one trillion tokens in about five hours.
  • Computationally efficient hash functions. We introduce novel non-cryptographic hash functions that reduce computational cost by partially reusing the previous result during deduplication. As a result, FED can process the hash generation part more than 1,700 times faster than the existing CPU baseline and over 100 times faster than the GPU baseline.
  • Stable deduplication quality. By comparing the deduplication results of the widely used MinHash implementation with our method, we have confirmed that FED provides both speed and reliability in deduplication.
  • Publicly available code base. We make the FED implementation publicly available. As LLMs continue to scale, FED will be instrumental in maintaining high data quality and enabling faster training cycles, ultimately advancing the field of NLP.
RealNews Dataset index: 12347 text: Margins matter. The more Synovis Life Technologies (Nasdaq: SYNO) keeps of each buck it earns in revenue, the more money it has to invest in growth, fund new strategic plans, or (gasp!) distribute to shareholders. Healthy margins often separate pretenders from the best stocks in the market. That’s why we check up on margins at least once a quarter in this series. …(more) index: 39083 text: Margins matter. The more Helen of Troy (Nasdaq: HELE) keeps of each buck it earns in revenue, the more money it has to invest in growth, fund new strategic plans, or (gasp!) distribute to shareholders. Healthy margins often separate pretenders from the best stocks in the market. That’s why we check up on margins at least once a quarter in this series. … (more)
Figure 1: Examples of duplicate documents.
2 Background and Related Work
2.1 Effects of Data Deduplication
Data deduplication involves identifying and removing redundant data to ensure each instance is unique, enhancing learning efficiency and improving data quality. Figure 1 illustrates duplicate data within the RealNews Zellers et al. (2019) dataset. The sentence structures of the two documents are nearly identical, and such redundancy can hinder effective learning for language models (LMs). Lee et al. (2022) suggests that dataset deduplication positively impacts the validation perplexity of language models, and Allamanis (2019) presents that duplicated data in code datasets negatively affects LMs on code understanding tasks. Moreover, train-test leakage, where test data is inadvertently included in the training set, can significantly skew model evaluation metrics. For example, near-duplicates in both training and test sets can lead to overestimating model performance. Thus, deduplication helps provide a more accurate assessment of model capabilities by eliminating these overlaps Lee et al. (2022).
Training on deduplicated datasets also reduces computational resource requirements, as the model processes unique information only once. This results in faster training times and lower costs. Deduplication also mitigates the risk of overfitting by preventing the model from memorizing repeated data entries, leading to better generalization of unseen data. Lee et al. (2022) states that deduplication also helps reduce some privacy concerns as it reduces the likelihood of models generating memorized content from training data, which could include sensitive information.
2.2 MinHash
MinHash Broder (1997) is a technique approximating the Jaccard similarity (please see Appendix A) between two documents. Identifying similar (duplicate) sets using MinHash can be broken down into four steps: MinHash generation, duplicate pair generation, constructing a union graph, and generating the final list of duplicates.
MinHash generation.
First, each document is broken down into smaller components, known as shingles or n-grams and represented by the set of its shingles. Then, the shingles are mapped to zero or positive integers using $H$ hash functions, $f_{1},f_{2},\cdots,f_{H}$. As a result, we have $H$ sequences of integers for a document. A MinHash signature for the document is a sequence of the minimum values obtained from each hash function, resulting in a sequence of length $H$. The Jaccard similarity between the two documents is then estimated by comparing their MinHash signatures. $H$ different hash functions result in a $H\times N$ signature matrix for $N$ documents. The choice of hash functions and the number of hash functions are critical hyperparameters in document deduplication based on MinHash. Please, see Appendix B for the details of MinHash generation.
Duplicate pair generation.
MinHash compares every pair of document signatures to determine if the two documents are similar (duplicates). MinHash approximates two documents $D_{1}$ and $D_{2}$’s similarity $sim_{doc}(D_{1},D_{2})$ with their signatures $S_{1}$ and $S_{2}$’s similarity $sim_{sig}(S_{1},S_{2})$:
$sim_{doc}(D_{1},D_{2})\approx sim_{sig}(S_{1},S_{2}).$ (1)
Specifically, MinHash approximates the Jaccard similarity by counting the same elements in the two signatures element-wise:
$sim_{sig}=\frac{\mbox{The \# of the same elements}}{\mbox{The \# of elements in a signature}}$ (2)
To classify a pair of documents as duplicates, MinHash uses a similarity threshold value $\theta$. If $sim_{sig}(S_{1},S_{2})>\theta$, the documents $D_{1}$ and $D_{2}$ are considered duplicates (i.e., similar).
Constructing a union graph.
After generating the list of duplicate pairs, we construct a union graph that clusters the duplicate pairs into groups. In the union graph, each document is a node, and an edge exists between two nodes if they form a duplicate pair. The resulting graph is an acyclic undirected graph where each connected component represents a group of near-duplicate documents.
Generating the final list of duplicates.
Based on the connected components identified in the previous step, we select a representative document (e.g., the document with the lowest document index) for each connected component to build the deduplicated dataset.
2.3 MinHash LSH
Since MinHash computes the similarity for every pair of documents, it leads to a significant computational cost. To solve this problem, MinHash LSH combines MinHash with Locality-Sensitive Hashing (LSH) (Indyk and Motwani, 1998).
The critical difference between MinHash and MinHash LSH lies in the stage of generating duplicate pairs. In MinHash LSH, the signature column vector of each document is divided into $b$ bands, each of which has $r$ integers. LSH hashes the bands in each document to $k$ buckets. When two documents have bands hashed to the same bucket, they are potentially similar documents. Please, see Appendix C for the details of MinHash LSH.
MinHash LSH is the most commonly used method for document deduplication Albalak et al. (2024). Brown et al. (2020) use MinHash LSH with ten hash functions to filter the training dataset for GPT-3. Rae et al. (2021) use MinHash LSH with 13-grams and 450 hash functions for the gopher training dataset.
Despite MinHash LSH is an approximation of MinHash, it poses a significant computational and execution time burden. Surprisingly, previous studies have not seriously considered the time required for the deduplication process.
2.4 Existing MinHash LSH Implementations
The MinHash-LSH algorithm, implemented using the Python library datasketch, is widely used in many NLP studies. SlimPajama Shen et al. (2023) employed CPU-based MinHash LSH using this library. We identified and fixed some bugs in their implementation, and hereafter, we will refer to this corrected version as the CPU baseline throughout the paper. The CPU baseline operates identically to the MinHash LSH process described in Section 2.3 and Appendix C, with one key difference.
The difference lies in how comparisons are performed between documents grouped into the same bucket. When a bucket contains $B$ documents, comparisons are performed with a $O(B^{2})$ time complexity. However, in the CPU baseline, comparisons are conducted between the first document’s signature vector entering the bucket and those of subsequently arriving documents, resulting in $O(B)$.
Recently, NVIDIA introduced a framework called NeMo Curator Jennings et al. , which includes a GPU-accelerated MinHash LSH algorithm using cuDF RAPIDS and Dask libraries Dask . We will refer to it as a GPU baseline. It is based on the approach used for Megatron-Turing NLG Smith et al. (2022). Similar to the CPU baseline, it does not compare all pairs of documents within a bucket containing $B$ documents. An anchor document for each bucket is selected, and the similarity between it and other documents in the bucket is computed. The GPU baseline is implemented using the Python GPU Data Frame library, cuDF, where document data is converted into a data frame format and processed with GPU-accelerated operations.
3 Design and Implementation of FED
We present a deduplication framework optimized for GPU cluster environments where each node has multiple GPUs. Considering scenarios where the size of the document set (e.g., trillions of tokens) exceeds the CPU and GPU memory capacity of the cluster, we adopt a strategy to store intermediate results, such as the signature matrix and bucket IDs, on storage devices during the deduplication process.
3.1 Minhash LSH
Assume that the dataset consists of multiple files and that they are stored in a suitable format, such as JSONL. Also, assume that each node in the cluster has $N_{\mbox{GPU}}$ processes when $N_{\mbox{GPU}}$ is the number of GPUs in each node. Each process reads and processes one file at a time. The file is loaded into the CPU’s main memory using two buffers according to a predefined format: one stores the text of all the documents in the file, and the other stores the document indices. These buffers are then transferred to the GPU. On the GPU, multiple CUDA threads generate the MinHash signature matrices for the documents in parallel and create bucket IDs for each band. The signature matrices and bucket IDs are transferred back to the CPU’s main memory and saved to a file. Consequently, each input file has a corresponding hash result file.
In MinHash LSH, documents sharing the same bucket ID are grouped together, and pairwise comparisons of their signature vectors are performed. Since documents with the same bucket ID are distributed across all hash result files, it requires reading all the hash result files, introducing a file I/O bottleneck. To address this problem, we store the documents with the same bucket ID in the same buffer and transfer the buffer to the GPU. Then, the GPU compares the document pairwise.
Unlike the CPU or GPU baseline implementation, which selects a representative document within a bucket containing $B$ documents and performs comparisons in $O(B)$, FED compares all pairs of documents in the bucket in $O(B^{2})$. Despite the increased number of comparisons, FED’s optimized GPU comparison kernel outperforms existing methods. By performing more comparisons, we aim to reduce false negatives.
The similarity computation results are then transferred back to the CPU, where similar documents are connected to build union graphs. Finally, these graphs are merged to group duplicate documents.
3.2 Hash Functions
MinHash LSH performs hashing twice. The first hashing involves calculating the MinHash values for each document to generate signature matrices. The second hashing assigns bucket IDs to the bands of the signature vector of each document. The CPU and GPU baselines and FED differ in implementing these hashing steps.
3.2.1 Hashing in the CPU baseline
The process of defining hash function $h_{i}$ for MinHash generation in the CPU baseline is described as follows:
  1. 1.
    Set a large Mersenne prime $p=2^{61}-1$.
  2. 2.
    Generate two sequences of random numbers $\{a_{i}\}_{i=1}^{H}$ and $\{b_{i}\}_{i=1}^{H}$ such that $0<a_{i}<p$ and $0\leq b_{i}<p$.
  3. 3.
    $h_{i}=(a_{i}\times\text{SHA-1}(u)+b_{i})\bmod p$,
where $H$ is the number of different hash functions, and $u$ refers to shingles composed of $n$-grams. CPU baseline uses a commonly used secure hash function SHA-1 Eastlake 3rd and Jones (2001), which is relatively slow when processing large data.
LSH divides the signature vector of each document into $b$ bands and hashes the values of each band into $k$ buckets. In the CPU baseline, documents are mapped to the same bucket only if their signature vectors are identical in a particular band. As a result, the bucket size, $k$, tends to be very large in the CPU baseline.
3.2.2 Hashing in the GPU Baseline
The primary focus of MinHash generation is preventing collisions rather than cryptographic security. However, the CPU baseline uses SHA-1, which guarantees cryptographic security. The GPU baseline uses the non-cryptographic MurmurHash3 algorithm for MinHash generation. It produces a 32-bit or 128-bit hash value, enabling faster computation compared to SHA-1.
In the LSH stage, the values of each band are mapped to buckets using the results of the MD5 hash function, which generates a 128-bit fingerprint by encoding a string of any length.
3.2.3 Hashing in FED
Unlike the CPU or GPU baseline, FED uses a computationally efficient and partially reusable non-cryptographic hash function for MinHash generation. Let $s=c_{1}c_{2}\ldots c_{k}$ be a $k$-gram shingle that consists of characters and $f(s)=\sum_{i=1}^{k}c_{i}\cdot q^{i-1}$ such that $q$ is a constant larger than the character alphabet size. Then FED’s hash function $h(s)$ for MinHash generation is defined as follows:
$h(s)=f(s)\bmod p$ (3)
where $p$ is a sufficiently large prime number. For example, we set $p=4294967$ in our experiment. In this case, both $p$ and $q$ can be represented with 32-bit integers. Unlike SHA-1 and MurMurHash3 in the CPU and GPU baselines, the proposed hash function in Equation 3 allows reusing the hash value of the previous shingle. For example, let $t=c_{2}\ldots c_{k}c_{k+1}$ be the next $k$-shingle of a shingle $s=c_{1}c_{2}\ldots c_{k}$. Then, $h(t)$ is computed as follows:
$\begin{array}[]{rl}h(t)&=f(t)\bmod p\\ &=\left(\frac{f(s)-c_{1}}{q}+c_{k+1}q^{k-1}\right)\bmod p\end{array}$ (4)
Instead of calculating $f(s)$ for each shingle, only a few operations – a multiplication, two additions, and a single division – are necessary when computing the hash value of the next shingle with that of the current shingle. This significantly reduces the computational cost. Moreover, unlike SHA-1, FED’s hash function is not constrained by the length of the string, making it more efficient for longer strings. We vary $(p,q)$ pairs to generate multiple hash functions simultaneously, supporting parallel computation. The hash function is suitable for MinHash with respect to the four requirements of good hash functions: determinacy, uniformity, collision resistance, and security. Security is not a concern because the hash function is not used for security purposes. Appendix D discusses the suitability of the hash functions for MinHash.
For hashing to generate bucket IDs, FED sums the $r$ values in each band and uses the remainder of the division by the predefined constant number of buckets, $K$. This effectively maps each band, which is defined in the space $\mathbb{N}^{r}$, into the space $\mathbb{N}^{K}$. As $K$ increases, the process of scanning through files to group documents with the same key becomes more frequent. The optimal number of buckets can be expressed as $K=kN^{1/2}$, where $k$ is a constant determined empirically (our experimental result shows that setting $k$ to a value between 1 and 2 is appropriate). Using this relationship, we identify the optimal value of $K$. Appendix E describes the details of obtaining $K$.
4 Optimizations and GPU Acceleration
This section explains the optimization details when calculating signature matrices and generating duplicate pairs.
Initial Processing by the CPU.
As mentioned in Section 3.1, each node in the target cluster has $N_{\mbox{GPU}}$ processes when $N_{\mbox{GPU}}$ is the number of GPUs in each node. The CPU processes handle different JSONL files in parallel. Each JSONL file is loaded into the CPU’s main memory using two buffers: one stores the text of all the documents in the file, and the other stores the document indices. After the buffers are filled with a batch of files, the buffer is transferred to the corresponding GPU for shingling and MinHash generation.
Shingling and MinHash generation on the GPU.
The GPU performs shingling and generates MinHash values for the documents contained in the buffer. Each JSONL file consists of multiple documents, and each SM (Streaming Multiprocessor) within the GPU processes different documents independently. That is, a thread block processes a document, with each thread handling different MinHash functions.
Communication-computation overlapping.
To maximize performance, we implement communication-computation overlapping techniques, such as double buffering. When the GPU performs computation, a buffer containing a batch of files is simultaneously transferred to the GPU.
Hash parameters for GPU compatibility.
The values for $p$ and $q$ described in Section 3.2.3 are carefully selected to address the limited support of the GPU’s 64-bit integer operations, ensuring that the result of the operation remains within the 32-bit integer range.
Storing back hash results and bucket IDs.
Before the pairwise comparison step of signatures, we transfer the hash results and bucket IDs to the CPU. If the CPU memory is insufficient, we store them in secondary storage, such as SSDs and HDDs, to address the issues in the target environment with a limited main memory space.
File I/O operation optimization.
To perform pairwise comparisons, we need to gather all documents that belong to the same bucket. That is, we must search the stored files to locate all documents that belong to a given bucket. Since performing file I/O for each bucket ID would be time-consuming, we search multiple bucket IDs within a single file I/O operation, resulting in a significant file I/O time reduction. The number of bucket IDs processed per file I/O operation is a tunable parameter. However, finding a suitable value that balances the CPU memory limits while sufficiently reducing the file I/O time is crucial.
GPU-based pairwise comparison.
Given $c$ documents in a bucket, each document has a signature vector $S_{i}$ of the length equivalent to the number of hash functions used, where $i\in\{1,\dots,c\}$. For any pair $(S_{i},S_{j})$, we need to compute the following:
$sim_{sig}(S_{i},S_{j})=\sum_{k}(S_{i,k}==S_{j,k}),$ (5)
where ==absent$=$ = is the equality operator.
Since typical libraries do not support $sim_{sig}(S_{i},S_{j})$, we implement a custom GPU kernel. Replacing the equality operator (==absent$=$ =) with the multiplication operator ($\times$) makes the operation matrix multiplication. Thus, the GPU kernel performs $sim_{sig}(S_{i},S_{j})$ in a matrix multiplication-like manner. The kernel exploits tiling and shared memory Harris (2012) to optimize memory access patterns and enhance parallel processing performance. After comparing $c$ signature vectors pairwise to calculate the Jaccard similarity, the index information of pairs with a Jaccard similarity higher than the predefined threshold is sent to the CPU.
5 Experiment
This section compares FED against the existing CPU and GPU implementations of MinHash LSH.
Table 1: System configuration
CPU 2 $\times$ AMD EPYC 7502 32-Core Processor Memory 8 $\times$ 64GB DDR4 DIMM GPU 4 $\times$ NVIDIA Tesla V100 OS Ubuntu 20.04.6 (kernel 5.4.0-100) Compiler nvcc 12.4, GCC 9.4 GPU Driver 520.61.05 CUDA Version 12.4
5.1 Target System Configuration
We use a 4-node GPU cluster for our experiments. Each node is equipped with four NVIDIA Tesla V100 GPUs, with each GPU having 32GB of memory. Table  1 provides the detailed target system configuration.
Table 2: The deduplication time in seconds on a single node is compared across the CPU baseline, the GPU baseline, and FED. The CPU baseline uses 64 CPU cores. Both the GPU baseline and FED are evaluated using four GPUs.
Method Dataset Generation Comparison Union Total Speedup
CPU baseline RealNews 18,361.1 2,983.7 17.6 21,361.4 1.0
(SlimPajama Shen et al. (2023)) C4 22,748.4 3,878.8 33.1 26,660.3 1.0
GPU baseline RealNews 1,102.4 2,016.1 36.1 3,154.7 6.8
(NVIDIA NeMo Curator Jennings et al. ) C4 1,340.8 2,992.6 31.1 4,364.5 6.1
FED RealNews 10.5 354.5 1.8 366.7 58.3
C4 13.1 696.3 1.9 711.4 37.5
5.2 Comparison Baselines
As mentioned in Section 2.4, the CPU baseline is the CPU-based MinHash LSH implementation found in SlimPajama Shen et al. (2023). We identified and fixed some bugs in the implementation. The GPU baseline is the GPU-accelerated MinHash LSH implementation in NVIDIA NeMo Curator Jennings et al. .
5.3 Datasets
The datasets we used in our experiments are the RealNews dataset Zellers et al. (2019) and C4 Raffel et al. (2020). The RealNews dataset is a large English corpus of news articles from Common Crawl, and C4 is a filtered version of Common Crawl. For C4, we sampled 100GB out of the total 700GB. We choose these datasets for two reasons. They are large-scale datasets containing over 30 million documents. We select a sufficiently large dataset to understand how long the traditional MinHash LSH method takes to process it and to see FED achieves how much acceleration. It is known from a previous study Lee et al. (2022) that deduplication of the datasets we choose can improve the performance of language models.
Preprocessing the datasets.
Following SlimPajama Shen et al. (2023), we preprocess the datasets before deduplication. We apply NFC normalization to remove non-Unicode characters, ensuring that a letter followed by a combining character becomes a single combined character. We also filter out documents with less than 200 characters.
For MinHash LSH, we use the default settings of the CPU baseline, 128 hash functions with $b=16$ and $r=8$. Following the approach by Lee et al. Lee et al. (2022), we create shingles using five-grams and set the Jaccard similarity threshold to 0.8. The default settings of the CPU and GPU baselines are used for parameters other than the above-mentioned. As mentioned in Section 3.2.3, the number of buckets $K$ is obtained on the fly based on the number of documents $N$, such that $K=2N^{1/2}$.
5.4 Processing Speed
We measure three different types of execution time for deduplication speed comparison: generation, comparison, and union times. The generation time involves computing the MinHash signature matrices and assigning bucket IDs for each band. The comparison time involves finding documents with the same bucket ID and calculating the Jaccard similarity between them. Finally, the union time involves constructing a union graph structure for all documents and ultimately identifying the documents to be removed.
We use RealNews and C4, which contain 32M and 50M documents, respectively. The CPU baseline processing time is measured using Python’s multiprocessing module across 64 CPU cores, while both the GPU baseline and FED are measured using four V100 GPUs. The results are summarized in Table 2.
For deduplication on the RealNews dataset, FED is about 58.3 times faster than the CPU baseline and about 8.6 times faster than the GPU baseline. Notably, due to the introduction of reusable hash functions, the MinHash generation phase alone achieves roughly 1,800 times faster than the CPU baseline and 110 times faster than the GPU baseline. Similarly, on the C4 dataset, FED achieves a speedup of 37.5 over the CPU baseline and about 6.1 over the GPU baseline.
The improvement in the comparison stage is less pronounced than in the generation stage. As described in Section 3.1, FED performs more actual comparisons than others. While it is possible to reduce the number of comparisons by following the method used in the CPU or GPU baseline, the FED implementation prioritizes minimizing false negatives by performing more comparisons. The result shows that performing more comparisons is feasible because optimized GPU kernels enable fast computation.
The processing speed for all methods is faster on the RealNews dataset than on the C4 dataset. It is attributed to the distinct characteristics of the two datasets. For both datasets, we use 100GB of data; however, RealNews consists of 32M documents with an average length of 3,335.17 characters (including spaces), whereas C4 consists of 50M documents with an average length of 2,191.15 characters (including spaces). Consequently, the higher number of documents in C4 increases the time taken for the generation and comparison phases.
Table 3: Comparing the deduplication accuracy of the MinHash LSH-based approaches and the standard MinHash algorithm. The Ratio column represents the total number of documents that have near-duplicate pairs divided by the total number of documents. Jaccard column shows the Jaccard similarity between the sets of all near duplicates by MinHash and each method.
(a) 0.1M documents
Ratio Jaccard Time
CPU MinHash 7,317 / 0.1M - 12,800.69 s
CPU baseline 7,284 / 0.1M 0.995 440.19 s
GPU baseline 7,144 / 0.1M 0.953 172.05 s
FED 7,175 / 0.1M 0.956 1.45 s
(b) 1M documents
Ratio Jaccard Time
CPU MinHash 289,229 / 1M - 9 days
CPU baseline 288,683 / 1M 0.998 1,326.03 s
GPU baseline 285,134 / 1M 0.976 522.12 s
FED 283,248 / 1M 0.966 4.41 s
5.5 Deduplication Accuracy
We verify if FED correctly identifies near-duplicate documents. Assuming that the result of MinHash, which performs pairwise comparisons on all document pairs, is the most accurate, we compare the indices of duplicated documents identified by the MinHash algorithm with those found by the CPU and GPU baselines and FED.
We use a 0.1M and 1M dataset sampled from RealNews for this experiment because the MinHash algorithm is time-consuming on large datasets. Our sampling is not entirely random to ensure many duplicated pairs even within the two smaller datasets. We first use MinHash LSH to identify similar groups, then select a proportion of each group to construct the subset.
In the deduplication process, duplicate pairs are identified and used to construct a union graph, from which groups of near-duplicate documents are formed. Each group consists of indices of similar (near-duplicate) documents. We define the set of all document indices with at least one near duplicate as the set of all near duplicates.
We evaluate the accuracy of deduplication methods by measuring the Jaccard similarity between the set of all near duplicates obtained using MinHash and that obtained using each method. A higher similarity indicates that the method identifies duplicate pairs more similarly to MinHash.
Table 3 compares the deduplication accuracy of each method. FED achieves a Jaccard similarity of 0.95 or higher to the CPU MinHash result. A similarity score of 0.95 or higher means that the method has identified near-duplicate documents almost identically. In fact, on the 1M dataset, out of the 283,248 documents identified as having duplicate pairs by FED, 281,355 overlapped with those identified by MinHash. This indicates that our method not only achieves faster deduplication than other methods but also identifies near-duplicate documents similar to the accurate CPU MinHash.
Table 4: Scalability with the number of GPUs (in seconds)
# of GPUs 1 2 4 8 16
Generation 51.78 24.86 10.47 7.31 3.58
Comparison 1,098.48 606.25 354.45 228.66 141.47
Union 0.89 1.36 1.81 1.99 1.65
Total Time 1,151.15 632.47 366.73 237.96 146.70
Table 5: Scalability with the dataset size (in seconds) using 16 GPUs
Dataset Generation Comparison Union Total
RealNews
(30B tokens)
3.58 141.47 1.65 146.70
RedPajama-1T
(1.2T tokens)
1016.01 17463.6 37.83 18,517.44
Table 6: Scalability with the number of hash functions in FED on the 0.1M RealNews dataset
# of hash
functions
# of
bands
# of rows
per band
Time (s) Jaccard
128 16 8 1.45 0.956
256 16 16 1.65 0.924
512 32 16 3.70 0.943
5.6 Scalability
By varying the number of GPUs, the size of the dataset, and the number of hash functions used in MinHash generation, we evaluate the scalability of FED.
Number of GPUs.
Table 6 shows the results of measuring execution time when varying the number of GPUs. Note that eight or more GPUs involve the multi-node system configuration. We use the RealNews dataset, which has 32 million documents. Even with just a single GPU, the method is nearly 20 times faster than the CPU baseline, and with 16 GPUs, the deduplication process completes in less than 147 seconds. This experiment demonstrated that our method works effectively in multi-node environments, and with more available node resources, even faster processing is possible.
Dataset size.
We increase the dataset size to verify whether FED can process trillions of tokens at a practical processing speed. The results of deduplication using 16 GPUs on the RealNews and RedPajama-1T Computer (2023) datasets are shown in Table 6. Recent LLMs are typically trained with more than trillions of tokens. The results demonstrate that FED can process trillions of tokens in just 5.1 hours.
Number of hash functions.
The number of hash functions used in the MinHash generation step is a key parameter that could impact the deduplication performance. For example, Rae et al. (2021) use 450 hash functions. We vary the number of hash functions from 128 to 256 and 512, and Table 6 shows the results. Despite changing the number of hash functions, FED maintained a Jaccard similarity of 0.92 or higher to the MinHash results, demonstrating that FED remains robust even with variations in the number of hash functions.
6 Conclusion
As the amount of data used for training LLMs continues to grow, the time required for data preprocessing has become increasingly significant. We present a framework, called FED, which performs MinHash LSH-based deduplication efficiently on GPUs. FED significantly outperforms the CPU-based deduplication tool included in SlimPajama by up to 58.3 times and the GPU-based deduplication tool included in NVIDIA NeMo Curator by up to 8.6 times with a node of four GPUs. Deduplication of 30 billion tokens by FED takes about 147 seconds and 1.2 trillion tokens takes about five hours in a multi-node and multi-GPU environment. Similar documents identified by FED show a Jaccard similarity of 0.95 or higher with those found using the original MinHash algorithm, indicating that FED achieves fast and accurate deduplication. Various experimental results highlight the practicality of FED for deduplication in large-scale datasets used for LLMs. We plan to make FED’s source code publicly available.
Limitations
As described in Section 3.1, unlike the baseline methods, FED performs comparisons for all pairs within a single bucket. While this strategy aims to reduce false negatives, it may result in excessive computations. Baseline methods achieve sufficient deduplication performance by selecting a representative document within each bucket and comparing it with the remaining ones. Thus, sensitivity experiments are required to determine how many comparisons are necessary for FED. If the same deduplication quality can be achieved with fewer comparisons, it would make FED faster. Detailed analysis and implementation are our future work.
Ethics Statement
We are committed to advancing NLP while adhering to ethical guidelines. Our study ensures compliance with data use policies, relying solely on publicly available datasets and open-source code, and strictly observing their licenses. Transparency is maintained through detailed descriptions of our methodologies, data sources, experimental processes, and the tools employed. Furthermore, we abstain from handling personal data or content that could perpetuate harm or discrimination against individuals or communities.
Acknowledgement
This work was supported in part by the National Research Foundation of Korea (NRF) grant (No. RS-2023-00222663, Center for Optimizing Hyperscale AI Models and Platforms), by the BK21 Plus programs for Innovative Data Science Talent Education Program (Dept. of Data Science, SNU, No. 5199990914569) through NRF, by the Institute for Information and Communications Technology Promotion (IITP) grant (No. 2018-0-00581, CUDA Programming Environment for FPGA Clusters), all funded by the Ministry of Science and ICT (MSIT) of Korea. ICT at Seoul National University provided research facilities for this study.
Appendix
Figure 2: The process of MinHash generation.
Figure 3: Hashing by MinHash LSH.
Appendix A Jaccard Similarity
Jaccard similarity (Jaccard, 1912), $J(A,B)$, is a metric to quantify the similarity of two finite sets, $A$ and $B$:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (6)
When used as a metric for finding document similarity, it is defined as the number of common words between the documents divided by the total number of words in them. While this is the most intuitive way to measure similarity, calculating the Jaccard similarity directly between documents in a large corpus is not feasible due to the computational cost. Thus, the MinHash algorithm Broder (1997), which approximates Jaccard similarity, is often used instead.
Appendix B MinHash Generation
Consider Figure 2 that illustrates the MinHash process. First, each document is broken down into smaller components, known as shingles or n-grams and represented by the set of its shingles. The shingles are typically sequences of contiguous words or characters. For example, the sentence "abcab" has a sequence of three shingles, $abc,bca,cab$. Figure 2 uses 3-grams for singling.
Then, we map the shingles to zero or positive integers using $H$ hash functions, $f_{1},f_{2},\cdots,f_{H}$. Each hash function assigns a unique number to a shingle, representing the document by a sequence of integers. As a result, we have $H$ sequences of integers for a document.
A MinHash signature for the document is a sequence of the minimum values obtained from each hash function, resulting in a sequence of length $H$, where $H$ is the total number of hash functions used. The Jaccard similarity between the two documents is then estimated by comparing their MinHash signatures. $H$ different hash functions result in a $H\times N$ signature matrix for $N$ documents, where each column represents the minimum hash values obtained from $H$ different hash functions.
For example, in Figure 2, the column $2$ in the signature matrix is the signature of document $D_{2}$. For $D_{2}$, the minimum value of the hash values by the hash function $f_{1}$ is $M_{{D_{2}},{f_{1}}}$, by $f_{2}$ is $M_{{D_{2}},{f_{2}}}$, and so on. The MinHash signatures are much smaller than the original sets of shingles, allowing efficient computation of similarity between documents. The choice of hash functions and the number of hash functions are critical hyperparameters in document deduplication based on MinHash.
Appendix C MinHash LSH
The critical difference between MinHash and MinHash LSH lies in the stage of generating duplicate pairs. In MinHash LSH, the signature column vector of each document is divided into $b$ bands, each of which has $r$ integers, as shown in Figure 3. LSH hashes the bands in each document to $k$ buckets. Let $S_{i}$ be the signature vector of the document $D_{i}$. For a given signature vector $S_{1}$ and $S_{2}$, when at least a pair of bands $B_{S_{1}}$ from $S_{1}$ and $B_{S_{2}}$ from $S_{2}$ hashed to the same bucket, we tag them as candidate pairs, potentially similar documents. For example, in Figure 3, $D_{2}$ and $D_{N}$ are potentially similar documents because they have bands hashed to the same bucket. For the candidate pairs, LSH performs pairwise comparisons on each document’s signature vectors of length $b\times r$. If the similarity between the two vectors exceeds a predefined threshold, the two documents are considered a duplicate pair. Once the union graph is constructed for each band, the union graphs for all bands are merged to obtain the final union graph.
Appendix D Suitability of FED’s Hash Functions
In general, a good hash function should have the following properties:
  • Determinacy: The same input produces the same output.
  • Uniformity: The hash function should distribute outputs uniformly across the range.
  • Collision resistance: Two distinct inputs should unlikely produce the same hash value.
The proposed hash function satisfies these properties. Determinacy and uniformity are straightforward. Collision resistance is also met with a large value of $p$. When uniformity is satisfied, the probability of a collision approaches $\frac{1}{p}$, which is very small. In many contexts, hash functions are required to resist preimage attacks, making it computationally difficult to reverse-engineer the input from the hash value. However, it is not a concern because the hash function is not used for security purposes.
Appendix E Number of Buckets in LSH
For hashing to generate bucket IDs, FED sums the $r$ values in each band and uses the remainder of the division by the predefined constant number of buckets, $K$. This effectively maps each band, which is defined in the space $\mathbb{N}^{r}$, into the space $\mathbb{N}^{K}$. As $K$ increases, the process of scanning through files to group documents with the same key becomes more frequent. This process is proportional to $K$, resulting in $O(K)$. On the other hand, the number of documents with the same key is, on average, $O\left(\frac{N}{K}\right)$ when $N$ is the total number of documents. Since the comparison time is proportional to the square of the number of documents, the overall time complexity becomes $O\left(\left(\frac{N}{K}\right)^{2}K\right)=O\left(\frac{N^{2}}{K}\right)$. Thus, the processing time can be approximated as $T(K)=aNK+\frac{bN^{2}}{K}$, where $a$ and $b$ are some constants. To minimize $T(K)$, we differentiate it with respect to $K$ and set $\frac{dT(K)}{dK}=0$, yielding $K^{2}=\frac{bN^{2}}{aN}$. Solving for $K$, we find $K=\sqrt{\frac{b}{a}}N^{1/2}$. Thus, the optimal number of buckets can be expressed as $K=kN^{1/2}$, where $k=\sqrt{\frac{b}{a}}$ is a constant determined empirically (our experimental result shows that setting $k$ to a value between 1 and 2 is appropriate). Using this relationship, we identify the optimal value of $K$.